OpenStudy (anonymous):

Complete the table of values for the following equation. 6x + 2y = 30 x-values y-values -1 0 15 1 2 9 3 6 4 Select the best answer from the choices provided. x-values y-values -1 18 0 15 1 12 2 9 3 6 4 2 x-values y-values -1 18 0 15 1 11 2 9 3 6 4 3 x-values y-values -1 17 0 15 1 12 2 9 3 6 4 2 x-values y-values -1 18 0 15 1 12 2 9 3 6 4 3

OpenStudy (texaschic101):

this is just a matter of subbing in numbers.. 6x + 2y = 30 .....if x = -1.....so sub in -1 for x to find y 6(-1) + 2y = 30 -6 + 2y = 30 2y = 30 + 6 2y = 36 y = 36/2 y = 18 so if x = -1, then y = 18 6x + 2y = 30 ....if x = 1 6(1) + 2y = 30 6 + 2y = 30 2y = 30 - 6 2y = 24 y = 24/2 y = 12 so if x = 1, then y = 12 I think you can get it now....just sub in the numbers given into the equation to find the other number.

OpenStudy (alex_mattucci):

It is the last one! D.